Give an example of a function of two variables whose first order partial derivatives exist at


(0, 0) (that is fx(0, 0) and fy(0, 0) both exist), but f is NOT differentiable at (0, 0).


Explain also why the function is NOT differentiable at (0, 0).

1
Expert's answer
2021-11-24T14:30:12-0500

Example

(a)Here is discontinuous function at (0,0) having partial derivatives at (0,0)


"f(x,y)= \begin{cases}
 \frac{2xy}{x^2+y^2} &\text{(x,y)} \neq (0,0)\\
 0 &\text{(x,y)} =(0,0)
\end{cases}"

showing f is not continuous at (0,0)

choosing the path x = 0 we see that "f(0,y)=0," so "lim_{y\to 0}f(0,y)=0".

choosing the path "x=y" we see that "f(x,x)=\frac{2x^2}{2x^2}=1," so "lim_{x\to 0}f(x,x)=1". The two path theorem implies that "lim_{(x,y)\to(0, 0)}f(x,y)" does not exist


(b) Finding "f_x(0,0) \ and\ f_y(0,0)"

the partial derivatives are defined at (0,0)

"f_x(0,0)=lim_{h\to 0}\frac{1}{h}[f(0+h,0)-f(0,0)]=lim_{h\to 0}\frac{1}{h}[0-0]=0"


"f_y(0,0)=lim_{h\to 0}\frac{1}{h}[f(0,0+h)-f(0,0)]=lim_{h\to 0}\frac{1}{h}[0-0]=0"


Therefore "f_x(0,0)=f_y(0,0)=0"